# M4-SAR: A Multi-Resolution, Multi-Polarization, Multi-Scene, Multi-Source Dataset and Benchmark for Optical-SAR Fusion Object Detection

### v1

@@ 2 M4-SAR Dataset
Data Annotation and Label Mapping. As the acquisition times of Sentinel-1 [50] and Sentinel-2 [9] are not perfectly synchronized, instance-level alignment of dynamic targets (e.g., ships) becomes difficult, which compromises the accuracy of label transfer. To address this issue, we focus on static and semantically meaningful targets, such as bridges, harbors, oil tanks, playgrounds, airports, and wind turbines, and annotate them using oriented bounding boxes. SAR’s unique imaging characteristics cause ambiguous target boundaries, challenging accurate annotation by non-experts. In contrast, optical images provide clearer visual cues that facilitate both target identification and annotation transfer. With precise alignment, annotations derived from optical images can be reliably transferred on SAR images, thereby significantly reducing manual annotation costs.
Given the scale of the dataset, fully manual labeling is both time-consuming and labor-intensive. To address this challenge, we propose a semi-supervised annotation strategy, as illustrated in Figure 2(c). The procedure involves the following steps: First, 5% of the entire dataset is manually annotated, ensuring that each selected image contains all six target categories: bridges, harbors, oil tanks, playgrounds, airports, and wind turbines. This subset establishes a high-quality foundation for subsequent annotation. Once manually labeled, a generic detector [21] is trained using this subset (with the same data used for both training and evaluation during the initial phase). The trained model is then employed to generate pseudo-labels for an additional 5% of the data, which are subsequently refined through manual correction and added to the labeled set. This updated set is used to retrain the detector in the next iteration. The annotation process is repeated iteratively until the full dataset is labeled. During this procedure, images without valid target instances are excluded to enhance annotation efficiency and the overall utility of the dataset.
Based on instance-level alignment, we achieve high-quality label transfer, enabling accurate projection of annotation information from optical images to their corresponding SAR counterparts. As a result, we construct M4-SAR, the first large-scale dataset for optical-SAR fusion object detection, which offers a solid foundation for subsequent method evaluation and algorithm development.
Statistical Analysis of Different Modalities. As shown in Table 1, our proposed M4-SAR dataset comprises 112,184 image pairs, 981,862 instances, and six representative target categories: bridges, harbors, oil tanks, playgrounds, airports, and wind turbines. Each image has a resolution of 512$\times$512 pixels and contains an average of 8.75 instances, highlighting the dense object distribution typical in remote sensing scenes. As illustrated in Figure 1 (c), rotated bounding boxes offer significantly higher localization accuracy compared to horizontal bounding boxes.
@@
  • Aspect Ratio: The aspect ratio of an object plays a critical role in detection accuracy. Figure 1 (b) illustrates the distribution of aspect ratios across categories. When the aspect ratio approaches 1, targets tend to appear square, potentially leading to recognition ambiguities. Conversely, extremely large aspect ratios increase the model’s sensitivity to object shapes, thus posing additional challenges to accurate detection.
@@
  • Angle: All target orientation angles are normalized to $\theta$ ($\theta\in[0,\pi/2)$), and their distributions are visualized in Figure 3(a). Notably, offshore wind turbines tend to exhibit square-like appearances, with orientation angles predominantly distributed around 0 and 90 due to the absence of surrounding reference structures and the top-down perspective of satellite imagery. Although oil tanks are circular, their labeled angles are contextually defined based on surrounding spatial features. This highlights the importance of contextual cues in remote sensing images and reflects the task’s sensitivity and complexity.
  • Number of Instances: Figure 3 (b) presents the instance distribution by category, where bridges and oil tanks constitute the majority of instances.
@@
  • Average Area of Objects: Figure 3 (c) shows the average pixel area per category, revealing the scale variability among different object categories in the dataset.
3 Benchmarks
@@
Optical-SAR Fusion Detection Tools (MSRODet). Due to the lack of publicly available methods specifically designed for optical-SAR fusion object detection, conducting a fair performance comparison remains challenging. To address this limitation, we propose an optical-SAR fusion object detection evaluation toolbox (MSRODet) built upon existing multi-source detection frameworks (e.g., RGB-infrared, RGB-depth). MSRODet integrates several representative methods, including CFT [42], CLANet [18], CSSA [2], CMADet [45], ICAFusion [45], and MMIDet [69], thereby providing a standardized and accessible benchmark for evaluating optical-SAR fusion object detection approaches. As shown in Figure 4 (a), empirical analysis reveals that significant domain gaps remain a primary obstacle to effective optical-SAR fusion. These discrepancies primarily arise from fundamental differences in visual characteristics and imaging mechanisms between optical and SAR modalities. Consequently, designing detection frameworks specifically tailored to optical-SAR fusion is essential for enhancing overall detection performance.
Figure 4: (a) Optical-SAR domain gap. (b) Overall framework of the proposed E2E-OSDet.
Figure 5: Architectural details of the proposed FAM, CMIM, and AFM modules.
@@
End-to-End Optical-SAR Fusion Object Detection (E2E-OSDet). Multi-source fusion object detection remains underexplored in the context of optical-SAR fusion, and significant limitations persist in addressing domain discrepancies between optical and SAR data. As shown in Figure 4 (b), to address these challenges, we propose a novel end-to-end optical-SAR fusion object detection framework, termed E2E-OSDet, which systematically addresses cross-modal challenges across three levels: data input, domain alignment, and feature fusion.
Filter Augment Module (FAM): First, to reduce the domain gap between optical and SAR images, we propose a Filter Augment Module (FAM) leveraging classical image filtering operators (e.g., HOG [5], Canny [1], Haar [52], Grad [24], and WST [40]). This module transforms the low-dimensional and sparse SAR representations into a more discriminative high-dimensional space, effectively mitigating the impact of domain differences.
Cross-modal Mamba Interaction Module (CMIM): Second, most existing methods focus on spatial-level fusion but often neglect the discrepancies in the feature domain. Mamba [14] possesses strong state-space modeling capabilities, which enables effective cross-modal sequence-level interactions and alleviates feature domain bias. However, current Mamba-based methods [70; 8] are hindered by complex architectures and intricate interaction mechanisms, limiting their practical applicability. To address this, we propose a Cross-modal Mamba Interaction Module (CMIM), which enables deep cross-modal interactions through simple and efficient operations.
Area-Attention Fusion Module (AFM): Finally, to further enhance the discriminability of fused features across different spatial regions, we introduce an Area-Attention Fusion Module (AFM). This module captures salient features of local targets and strengthens responses in critical regions, thereby enhancing the effectiveness of feature fusion. The detailed structure of each module is illustrated in Figure 5. For additional E2E-OSDet details, please refer to Appendix C.
4 Experiments
4.1 Setups and Implementation Details
Experiments were conducted on two NVIDIA RTX 4090 GPUs. All fusion models were trained over 300 epochs with the SGD optimizer, using a learning rate of 0.01, momentum of 0.937, and weight decay of 5$\times$10-4. The batch size was set to 64, and the input images were resized to a resolution of 512$\times$512. To ensure fair performance evaluation, all methods were evaluated using a unified framework comprising the YOLOv11 [21] backbone and the oriented bounding box (OBB) detection head from YOLOv8 [21]. All fusion models were trained from scratch, without leveraging any pre-trained weights. During evaluation, we adopted the standard COCO evaluation protocol to measure Average Precision (AP), including $AP_{50}$, $AP_{75}$, and mean AP ($mAP$), where $mAP$ denotes the mean AP across IoU thresholds ranging from 0.5 to 0.95. For more details, see Appendix D.
4.2 Quantitative Evaluation
Main Results. We compared multi-source fusion methods (e.g., CFT, CLANet, CSSA, CMADet, ICAFusion, MMIDet, and E2E-MFD) with single-source methods (e.g., RTMDet, PSC, KFIoU, LSKNet and YOLO series) that utilize either optical or SAR data exclusively. As shown in Table 2, the detection accuracy achieved using SAR data alone is considerably lower than that of the optical-only counterpart. This performance gap is primarily attributed to the unique imaging principles of SAR, which hinder the accurate estimation of bounding box shapes—even when targets are visually identifiable. All multi-source fusion methods outperform their corresponding unimodal baselines, demonstrating the effectiveness of cross-modal fusion in leveraging SAR data to complement optical information (and vice versa), while significantly reducing false positives. Notably, our proposed method, E2E-OSDet, achieves the highest performance among all fusion-based approaches, reaching an $mAP$ of 61.4%, which underscores the importance of mitigating domain discrepancies to improve detection accuracy. Furthermore, E2E-OSDet is also highly efficient, with only 27.5M parameters and an inference time of 20.9ms. Finally, we evaluate the generalizability of E2E-OSDet across various detection frameworks, with the results presented in Figure 6(e).
@@ 4.2 Quantitative Evaluation
Table 2: Quantitative results on the M4-SAR. O and S denote optical and SAR data, respectively. Inf.T indicates inference time per image. The best two results are highlighted in red and green.

### v2

@@ II-C Multi-source Fusion Object Detection
III M4-SAR Dataset
III-A Overview of the Proposed Dataset Construction Pipeline
Fig. 2 illustrates the overall pipeline of the proposed dataset construction process. In the first stage, we acquire optical and SAR images from Sentinel-1 and Sentinel-2 satellites provided by the European Space Agency (ESA). In the second stage, the acquired images are pre-processed and aligned. After completing the alignment, we crop aligned image pairs into fixed-size patches to generate initial unlabeled samples. In the third stage, we introduce a semi-supervised annotation strategy guided by optical images to ensure high-quality labeling of the image pairs.
III-B Data Collection Procedure
First, we collected the required data via the open-access platform provided by the ESA. A map-based selection tool was used to specify areas of interest, covering representative coastal cities, including Tianjin, Los Angeles, Hong Kong, and London. We then retrieved SAR images along with their corresponding digital elevation model (DEM) data according to specific criteria. For optical imagery, Sentinel-2 images covering the same regions as the Sentinel-1 data were acquired, specifically including the red (R), green (G), and blue (B) spectral bands. To minimize variations caused by temporal discrepancies between the two modalities, we selected image pairs with a time interval of less than 10 days. This setting ensures that targets of interest (e.g., bridges, oil tanks) maintain structural consistency within the short interval, thereby enhancing the accuracy of subsequent instance-level alignment. The product type for optical images is L2A, with cloud coverage limited to 0–10% (the corresponding cloud-free (0%) images are used for high-quality annotation) to guarantee high visual fidelity. The SAR images are provided in the Single Look Complex (SLC) format to preserve full coherence information. The collected data span the period from 2020 to 2022 and encompass multiple spatial resolutions and polarization modes. The optical images have 10m, 20m, and 60m resolutions, and the SAR data cover both vertical-horizontal (VH) and vertical-vertical (VV) polarizations.
III-C Data Processing
We combined the red, green, and blue channels to generate standard optical images using data processing software. The optical images were provided in three spatial resolutions: 10m, 20m, and 60m. To highlight the complementary characteristics of optical and SAR images, we retained only the high-resolution (10m) and low-resolution (60m) variants. For the SAR data, we first applied multi-look averaging to the SLC images in both the azimuth and range directions to reduce speckle effects and improve radiometric resolution. Subsequently, filtering techniques were employed to suppress speckle noise. Using the corresponding DEM data, we geocoded and radiometrically calibrated the SAR images and conducted geometric alignment and correction to improve spatial localization precision. After preprocessing, the optical and SAR images were aligned based on their geographic coordinates. As the overlap between the two modalities was partial, the aligned image pairs were cropped into $\times1,024$ patches, and only those with valid overlapping regions were retained for subsequent annotation. Notably, this initial cropping was intended to facilitate efficient manual annotation. After annotation was completed, we further partitioned the annotated images into non-overlapping $\times512$ tiles to meet the demands of different downstream applications.
@@ III-C Data Processing
III-D Data Annotation and Label Mapping
As the acquisition times of Sentinel-1 [44] and Sentinel-2 [7] are not perfectly synchronized, instance-level alignment of dynamic targets (e.g., ships) becomes difficult, which compromises the accuracy of label transfer. To address this issue, we focus on static and semantically meaningful targets, such as bridges, harbors, oil tanks, playgrounds, airports, and wind turbines, and annotate them using oriented bounding boxes. SAR’s unique imaging characteristics cause ambiguous target boundaries, challenging accurate annotation by non-experts. In contrast, optical images (cloudless) provide clearer visual cues that facilitate both target identification and annotation transfer. With precise alignment, annotations derived from optical images can be reliably transferred to SAR images, thereby significantly reducing manual annotation costs.
Given the scale of the dataset, fully manual labeling is both time-consuming and labor-intensive. To address this challenge, we propose a semi-supervised annotation strategy, as illustrated in Fig. 2 (c). The procedure involves the following steps: First, 5% of the entire cloudless optical images are manually annotated, ensuring that each selected image contains all six target categories: bridges, harbors, oil tanks, playgrounds, airports, and wind turbines. This subset establishes a high-quality foundation for subsequent annotation. Once manually labeled, a detector [16] is trained using this subset (with the same data used for both training and evaluation during the initial phase). The trained model is then employed to generate pseudo-labels for an additional 5% of the data, which are subsequently refined through manual correction and added to the labeled set. This updated set is used to retrain the detector in the next iteration. The annotation process is repeated iteratively until the full dataset is labeled. During this procedure, images without valid target instances are excluded to enhance annotation efficiency of the dataset.
During the annotation process, we use cloud-free images to ensure annotation accuracy. After annotation is complete, the cloud-free images are replaced with the corresponding cloud-containing images. Based on instance-level alignment, we achieve high-quality label transfer, enabling accurate projection of annotation information from optical images to their corresponding SAR counterparts. As a result, we construct M4-SAR, the large-scale dataset for optical-SAR fusion object detection, which offers a solid foundation for subsequent method evaluation and algorithm development.
III-E Statistical Analysis of Different Modalities
As shown in Table I, our proposed M4-SAR dataset comprises 112,174 image pairs, 981,862 instances, and six representative target categories: bridges, harbors, oil tanks, playgrounds, airports, and wind turbines. Each image has a resolution of $\times512$ pixels and contains an average of 8.75 instances, highlighting the dense object distribution typical in remote sensing scenes. As illustrated in Fig. 1 (c), rotated bounding boxes offer significantly higher localization accuracy compared to horizontal bounding boxes.
@@ III-E Statistical Analysis of Different Modalities
  • Aspect Ratio: The aspect ratio of an object plays a critical role in detection accuracy. Fig. 1 (b) illustrates the distribution of aspect ratios across categories. When the aspect ratio approaches 1, targets tend to appear square, potentially leading to recognition ambiguities. Conversely, extremely large aspect ratios increase the model’s sensitivity to object shapes, thus posing additional challenges to accurate detection.
@@
  • Angle: All target orientation angles are normalized to $\theta$ ($\theta\in[0,\pi/2)$), and their distributions are visualized in Fig. 3 (a). Notably, offshore wind turbines tend to exhibit square-like appearances, with orientation angles predominantly distributed around 0 and 90 due to the absence of surrounding reference structures and the top-down perspective of satellite imagery. Although oil tanks are circular, their labeled angles are contextually defined based on surrounding spatial features. This highlights the importance of contextual cues in remote sensing images and reflects the task’s sensitivity and complexity.
  • Number of Instances: Fig. 3 (b) displays the distribution of instances by category, revealing that bridges and oil tanks constitute the largest proportions in the dataset. This imbalance reflects the actual occurrence frequency of typical targets in coastal regions: in such areas, static infrastructure such as bridges and oil tanks dominates satellite imagery. The relatively low number of port instances is primarily due to their extensive spatial coverage, with a single port typically corresponding to a single annotated instance.
@@
  • Average Area of Objects: Fig. 3 (c) shows the average pixel area per category, revealing the scale variability among different object categories in the dataset.
Figure 4: (a) optical-SAR domain gap. (b) Overall framework of the proposed end-to-end optical-SAR fusion object detection (E2E-OSDet). Architectural details of the proposed filter augment module (c), cross-modal Mamba interaction module (d), and area-attention fusion module (e).
IV Method
IV-A optical-SAR Fusion Detection Tools
@@ IV Method
Due to the lack of publicly available methods specifically designed for optical-SAR fusion object detection, conducting a fair performance comparison remains challenging. To address this limitation, we propose a Multi-Source Rotated Detector (MSRODet) built upon existing multi-source detection frameworks (e.g., RGB-infrared, RGB-depth). MSRODet integrates several representative methods, including CFT [36], CLANet [14], CSSA [2], CMADet [39], ICAFusion [38], and MMIDet [60], thereby providing a standardized and accessible benchmark for evaluating optical-SAR fusion object detection approaches. As shown in Fig. 4 (a), empirical analysis reveals that significant domain gaps remain a primary obstacle to effective optical-SAR fusion. These discrepancies primarily arise from fundamental differences in visual characteristics and imaging mechanisms between optical and SAR modalities. Consequently, designing detection frameworks specifically tailored to optical-SAR fusion is essential for enhancing overall detection performance.
IV-B End-to-End optical-SAR Fusion Object Detection
IV-B1 Overall Framework
@@
Multi-source fusion for object detection remains relatively underexplored in the context of optical-SAR integration, and notable challenges persist due to the significant domain discrepancies between the two modalities. As illustrated in Fig. 4 (b), we propose an end-to-end optical-SAR fusion framework, E2E-OSDet, which explicitly tackles these cross-modal challenges at three complementary levels: data input, domain alignment, and feature fusion.
IV-B2 Filter Augment Module
To address the inherent domain disparities between optical and SAR images, caused by differences in imaging mechanisms and feature distributions, we propose the Filter Augment Module (FAM), a novel preprocessing framework designed to harmonize multi-source inputs for robust object detection. FAM incorporates a suite of classical image filtering operators, including HOG [4], Canny [1], Haar [45], Grad [21], and WST [34]. These filters extract complementary low-level cues, such as edges, gradients, and textures, under diverse imaging conditions. By projecting the sparse, low-dimensional features of both optical and SAR images into a high-dimensional, discriminative feature space, FAM reduces cross-modal domain gaps and enhances the alignment of modality-specific characteristics. Specifically, given an optical image $I_{O}\in\mathbb{R}^{H\times W\times 3}$ and a SAR image $I_{S}\in\mathbb{R}^{H\times W\times 3}$, FAM generates a set of enhanced representations $F_{O}\in\mathbb{R}^{H\times W\times 3}$ and $F_{S}\in\mathbb{R}^{H\times W\times 3}$ for each filter $t\in\{\text{WST},\text{Canny},\text{Haar},\text{HOG},\text{Grad}\}$, defined as:
$F_{O}=I_{O}+\alpha*\mathbb{FAM}^{t}(I_{O}),$ (1)
$F_{S}=I_{S}+\alpha*\mathbb{FAM}^{t}(I_{S}),$ (2)
$t\in\{\text{WST},\text{Canny},\text{Haar},\text{HOG},\text{Grad}\},$ (3)
where $\mathbb{FAM}^{t}(\cdot)$ denotes the $t$-th filter-specific augmentation function, and $\alpha$ represents a set of learnable scalars that balance the contributions of the original and filtered features for the optical and SAR modalities, respectively. This dual-branch enhancement scheme preserves the structural and textural integrity of each modality while injecting noise-robust, multi-scale edge and shape cues derived from classical filters. By aligning feature representations across modalities at the preprocessing stage, FAM facilitates more effective optical-SAR fusion. It addresses the limitations of conventional preprocessing pipelines and significantly enhances the generalization capability of multi-source object detection models under diverse and challenging remote sensing conditions.
Figure 5: The scanning mechanisms explored in this work include the Bidirectional Scan, Z-Order, Zigzag, and the 8-direction Hilbert Scanning Method.
IV-B3 Cross-modal Mamba Interaction Module
Existing multi-source fusion methods primarily emphasize spatial-level feature integration, often neglecting the substantial domain discrepancies between heterogeneous modalities such as optical and SAR images. These domain shifts pose a significant challenge to effective cross-modal fusion. The Mamba framework [10], with its state-space modeling and efficient sequence processing capabilities, has demonstrated strong potential for capturing long-range dependencies. This makes it well-suited for mitigating inter-modal inconsistencies. However, existing Mamba-based fusion methods [61, 6] typically involve intricate architectures and computationally intensive interaction designs, which limit their scalability and practical deployment. To overcome these limitations, we propose the Cross-modal Mamba Interaction Module (CMIM), a lightweight yet powerful module tailored to align optical and SAR features by leveraging Mamba’s long-sequence modeling strength. Given paired image features ($F_{Oj},F_{Sj}\in\mathbb{R}^{\frac{H}{2^{j}}\times\frac{W}{2^{j}}\times C_{j}},j\in\{3,4,5\}$), we first flatten them into one-dimensional sequences ($X_{j}\in\mathbb{R}^{1\times n}$, $Y_{j}\in\mathbb{R}^{1\times n}$) in row-major order:
$X_{j}=[p^{o}_{1},p^{o}_{2},\dots,p^{o}_{n-1},p^{o}_{n}],Y_{j}=[p^{s}_{1},p^{s}_{2},\dots,p^{s}_{n-1},p^{s}_{n}],$ (4)
where $p^{o}_{i}$ and $p^{s}_{i}$ represent patches in optical and SAR imagery, respectively. Traditional methods concatenate these sequences to obtain $Z^{tra}_{j}\in\mathbb{R}^{1\times 2n},j\in\{3,4,5\}$:
$Z^{tra}_{j}=[p^{o}_{1},p^{o}_{2},\dots,p^{o}_{n-1},p^{o}_{n},p^{s}_{1},p^{s}_{2},\dots,p^{s}_{n-1},p^{s}_{n}].$ (5)
However, these sequence modeling methods often fail to establish explicit local correspondences between modalities, leading to suboptimal cross-modal alignment and feature fusion.
To address this issue, we introduce an Interleaved Input Rearrangement (IIR) mechanism that enforces fine-grained spatial correspondence by interleaving paired features in a one-to-one fashion. Specifically, given vectorized feature sequences $X_{j}$ and $Y_{j}$ from the optical and SAR modalities, respectively, IIR constructs a unified sequence $Z^{IIR}_{j}\in\mathbb{R}^{1\times 2n}$ by interleaving the features as:
$Z^{IIR}_{j}=[p^{o}_{1},p^{s}_{1},p^{o}_{2},p^{s}_{2},p^{o}_{3},p^{s}_{3},\dots,p^{o}_{n-1},p^{s}_{n-1},p^{o}_{n},p^{s}_{n}].$ (6)
This interleaving ensures that corresponding positional features from both modalities are adjacent, enabling Mamba’s state-space modeling to directly capture semantic and structural alignments, thus enhancing domain coherence. The resulting interleaved sequence is subsequently fed into a directional scanning mechanism (see Fig. 5 [15]), which applies Mamba blocks along both horizontal and vertical directions. This bi-directional sequence modeling allows CMIM to capture long-range dependencies while preserving fine-grained local correspondence, facilitating deeper cross-modal interaction. Finally, the output sequence is reshaped back into the 2D spatial domain and fused with residual skip connections to retain original modality-specific information. Through this design, CMIM effectively bridges modality gaps and enhances the robustness of multi-source object detection under challenging remote sensing conditions.
$Z^{scan}_{j}=\mathbb{SCAN}^{m}(Z^{IIR}_{j}),$ (7)
$m\in\{\text{Bid, Z-Order, Zigzag, Hilbert}\}.$ (8)
$\{F^{\prime}_{Oj},F^{\prime}_{Sj}\}=\mathbb{CMIM}(Z^{scan}_{j}),\quad j\in\{3,4,5\}.$ (9)
@@ IV-B3 Cross-modal Mamba Interaction Module
TABLE II: Quantitative results on the M4-SAR. O and S denote optical and SAR data, respectively. Param. denotes the parameter count of the model. Inf.T indicates inference time per image. The best two results are highlighted in Black and mark.